Find the integral surface of π‘₯2𝑝 + 𝑦2π‘ž + 𝑧2 = 0 which passes through the curve x+y = xy, 𝑧= 1.


1
Expert's answer
2021-09-30T18:03:11-0400

x2p+y2q+z2=0

x2p+y2q=-z2

By lagrange equation;

"\frac{dx}{x^{2}}=\frac{dy}{y^{2}}=\frac{dz}{-z^{2}}"

Taking the 1st and 2nd terms,

"\frac{dx}{x^{2}}=\frac{dy}{y^{2}}"

Integrate both sides,

"\int \frac{dx}{x^{2}}=\int \frac{dy}{y^{2}}"

"C_1+\frac{1}{-x}=-\frac{1}{y}"

"C_1=\frac{1}{x}-\frac{1}{y}" This is equation (i)

Taking the 2nd and 3rd terms,

"\frac{dy}{y^{2}}=\frac{dz}{-z^{2}}"

Integrate both sides

"\int \frac{dy}{y^{2}}=\int \frac{dz}{-z^{2}}"

"-\frac{1}{y}+C_2=\frac{1}{z}"

"C_2=\frac{1}{y}+\frac{1}{z}" This is equation (ii)

The equation of the curve given is x+y=xy and z=1

We let x=t

Therefore from the equation of the curve;

x=xy-y

Rearrange; xy-y=x

y(x-1)=x

"y=\frac{x}{x-1}\implies y=\frac{t}{t-1}"

z=1

Put the above values of x,y and z in equations (i) and (ii) and remove t from the obtained relation

"C_1=\frac{1}{t}-\frac{t-1}{t}"

"C_1=\frac{1}{t}-1+\frac{1}{t}"

"C_1=\frac{1}{t}+\frac{1}{t}-1" This is equation (iii)

"C_2=1-\frac{1}{t}+1"

"C_2=2-\frac{1}{t}"

"\frac{1}{t}=2-C_2" Replace this in equation (iii)

"C_1=2-C_2+2-C_2-1"

"C_1=3-2C_2" This is equation (iv)

Now replace the values of C1 and C2 in (iv) to get the required surface

"\frac{1}{x}-\frac{1}{y}=3-\frac{2}{y}-\frac{2}{z}"

"\frac{y-x}{xy}=\frac{3yz-2z-2y}{yz}"

"zy-xz=3xyz-2xz-2xy"

The Integral surface is;

"zy+xz+2xy=3xyz"